how many liters of a 80% acid solution must be mixed with a 30% acid solution to get 300 liters of a 70% acid solution?
Answer by checkley79(3341)

.80X+.30(300-X)=300*.70
.80X+90-.30X=210
.50X=210-90
.50X=120
X=120/.50
X=240 LITERS OF 80% ARE USED.
300-240=60 LITERS OF 30% ARE USED.
PROOF:
.80*240+.30*60=210
192+18=210
210=210